You are excited to start investing and have found an ETF that you hope will continue to earn an interest rate of 10% per year. You plan to make bi-weekly deposits of $50 and you will make your first deposit today. How much will you have in 5 years? (Enter only numbers and decimals in your response. Round to 2 decimal places.)
Answer #1

Assuming bi weekly compounding

Rate = 10%/26

A= 50

N= 5*26 = 130

Future Value of annuity due= (1+Rate)*A*((1+rate)^number of periods-1)/rate

= (1+10%/26)*50*((1+10%/26)^130-1)/(10%/26)

=8445.19